OpenStudy (anonymous):

Two tuning forks are played at the same time. One has a frequency of 490 Hz and the other is 488 Hz. How many beats per second are heard? 490 488 2 4

OpenStudy (anonymous):

If the second harmonic of a certain string is 42 Hz, what is the fundamental frequency of the string? 21 Hz 42 Hz 84 Hz 126 Hz

OpenStudy (sidsiddhartha):

beats per second are heard=(490-488)=2 hz

OpenStudy (sidsiddhartha):

fundamental freq=42/2=21 hz